Provide the major organic product for each of the following transformations.


Provide the major organic product for each of the
Answer #1
Concepts and reason

The carbonyl compounds readily undergo a nucleophilic addition reaction that involves the breaking of carbon-oxygen (C=O)\left( {{\rm{C = O}}} \right) double bond. The reactivity of carbonyl compounds is due to the electronegative difference between oxygen and carbon. The oxygen attached to carbon is more electronegative. So, the oxygen can partially pull the electrons from the carbon atom making it more electron-deficient. Due to this, the nucleophile readily attacks the carbonyl carbon and undergoes an addition reaction.

Fundamentals

The reaction of carbonyl compounds is given below:

Wittig reaction: This method is used in the organic synthesis for the preparation of alkenes from carbonyl compounds. The carbonyl compounds on reaction with phosphorus ylides undergoes (2+2) a cycloaddition reaction and forms a four membered cyclic intermediate (oxophosphetane) followed by the formation of phosphine oxide to give alkenes.

The reaction scheme is given below:

o + (Phype_EM
CH2 + (Ph)3P=
0

Cyanohydrin formation: Carbonyl compounds on reaction with hydrogen cyanide (HCN)\left( {{\rm{HCN}}} \right) undergo a nucleophilic addition reaction to give cyanohydrin as a product.

The reaction scheme is given below:

HCN
R
R2

Reaction of amine derivatives with carbonyl compounds:

The carbonyl compounds on reaction with amines undergo an addition-elimination reaction to give imine derivatives.

The reaction scheme is given below:

berminata L
+ H2O
+ H2N-R
- H
R
CR,
Ri

Carbonyl reduction: The carbonyl compound on reaction with Lithium aluminium hydride (LiAlH4)\left( {{\rm{LiAl}}{{\rm{H}}_4}} \right) undergoes a reduction to give alcohols.

The reaction scheme is given below:

betw-
LiAlH4
Ri
R2

Protection of carbonyl compounds: Carbonyl compounds on reaction with 1,2-diols(protecting group) forms cyclic acetals.

OH
-
R2

The mechanism for the reaction of cyclohexanone with (HCN)\left( {{\rm{HCN}}} \right) is given below:

الا
Cyanohydrin product

The mechanism for the reaction of cyclohexanone with ethylene glycol is given below:

-Н,0
ДН,0
LOH
НО.
НО,
9
он
н

The mechanism for the reaction of cyclohexanone with phosphorus ylides is given below:

-P(Ph)3
P(Ph)3
CH₂
CH2
Ph
Ph
+
+
SP=o
Ph

The mechanism for the reaction of cyclohexanone with pyrrolidine is given below:

- HO
H
:
:
HO:

The product for the reaction of cyclohexanone with aniline is given below:

NH2
-*
primary amine
anotao
IZ
(8)
0H -
-Ze

The mechanism for the reduction of cyclohexanone is given below:

H
ZH
+
!1 /
30:
H
AIHz Lit

Ans:

The major organic product for each of the following transformation is given below:

HO
CN
1)LiAlH4
2) H
HCN
он
HO
trace H+
(CGHz)3P+_CH?
NH
trace H+